OpenStudy (anonymous):

How to factor 2x^2+5x-3???

jimthompson5910 (jim_thompson5910):

Have you learned the quadratic formula?

OpenStudy (anonymous):

Yes. Would this be applied here?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes the best way to factor in my opinion is to first find the roots using the quadratic formula. What roots do you get?

OpenStudy (anonymous):

25-(24)=1

jimthompson5910 (jim_thompson5910):

ok that looks like the value of b^2 - 4ac

jimthompson5910 (jim_thompson5910):

notice that since the discriminant is equal to 1, which is a positive perfect square, this means we can factor

OpenStudy (anonymous):

I see... so how do we proceed?

OpenStudy (anonymous):

I attempted to utilize the AC method when going about factoring the trinomial but ended up with an incorrect result.

jimthompson5910 (jim_thompson5910):

so do you have this written down? \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-5\pm\sqrt{5^2-4*2*(-3)}}{2*2}\] \[\Large x = \frac{-5\pm\sqrt{49}}{4}\] notice how 49 is under the root and NOT 1

OpenStudy (anonymous):

Yes; would we proceed by simplifying the division and the square root? -5+/-7/4= 1/2 or -3.

jimthompson5910 (jim_thompson5910):

correct, the two roots are 1/2 and -3

jimthompson5910 (jim_thompson5910):

if x = 1/2, then we can multiply both sides by 2 to get 2x = 1 then subtract the 1 from both sides to get 2x - 1 = 0 agreed?

OpenStudy (anonymous):

Certainly.

jimthompson5910 (jim_thompson5910):

then if x = -3, we can add 3 to both sides to get x+3 = 0

OpenStudy (anonymous):

I see where you're going now, by discovering the roots, it's much easier to find the factors.

jimthompson5910 (jim_thompson5910):

so the two roots x = 1/2 or x = -3 turns into these two equations 2x-1 = 0 or x+3 = 0

jimthompson5910 (jim_thompson5910):

yes, we then use the zero product property to go from `2x-1 = 0 or x+3 = 0` to `(2x-1)(x+3)=0`

jimthompson5910 (jim_thompson5910):

the last thing to do is to expand out `(2x-1)(x+3)` and see if you get `2x^2+5x-3`

OpenStudy (anonymous):

While this certainly is an easier method, I don't know if in application I will be able to know to multiply the root, x=1/2, by 2 on both sides. What tells us we need to do this?

jimthompson5910 (jim_thompson5910):

the goal is to get each root to turn into an equation where the right side is 0 so we can use the fact that `A = 0 or B = 0` turns into `A*B = 0`

jimthompson5910 (jim_thompson5910):

you dont' have to multiply by 2, but I don't like fractions which is why I try to get rid of them as soon as possible

jimthompson5910 (jim_thompson5910):

you could easily have x - (1/2) = 0 and x+3 = 0 turn into (x - (1/2))*(x+3) = 0 but that's a bit more complicated than it needs to be

OpenStudy (anonymous):

I guess I would remember to do it as the factors typically need to be written with whole numbers.

jimthompson5910 (jim_thompson5910):

yeah that's a good general rule to have so it's as simple as possible

OpenStudy (anonymous):

Thank you so much for your thorough response!

jimthompson5910 (jim_thompson5910):

you're welcome